OpenStudy (volleyballlover55):

What is the simplified expression for 4 power 4 multiplied by 4 power 3 over 4 power 5? 4^0 4^1 4^2 4^3

OpenStudy (volleyballlover55):

@CountryBoy15

OpenStudy (countryboy15):

hey

OpenStudy (countryboy15):

thats why okay what one = 16???

OpenStudy (countryboy15):

@volleyballlover55

OpenStudy (volleyballlover55):

idk

OpenStudy (countryboy15):

okay ill just tell u this time its = 4^2=16 and thats your answer okay

OpenStudy (volleyballlover55):

theres no option for that answer

OpenStudy (countryboy15):

also if i get baned its bc of juliusmatos12 he said hes spaming report on me so if im like ofline then you know what happend

OpenStudy (countryboy15):

what do you mean

OpenStudy (countryboy15):

the answer is 4^2 im just saying that 4^2=16 and 16 was the answer now you had to make it into a power so its 4^2 @volleyballlover55

OpenStudy (countryboy15):

okay need any more help let me know

OpenStudy (volleyballlover55):

ok

OpenStudy (countryboy15):

sorry i saw i wrought that incorrectly lol

OpenStudy (countryboy15):

medals

OpenStudy (countryboy15):

lol